# {cutout}03.2cm55cm1 empty PDF-WuKong         : A Large Multimodal Model for Efficient Long PDF Reading with End-to-End Sparse Sampling

### v2

@@ 3.1 Overview
Our pipeline consists of three components: a document parser, a sparse sampler, and a large language model, as shown in Fig. 2. The document parsing stage converts input PDFs into machine-readable content with interleaved text and images. The sparse sampler then encodes and caches embeddings for text blocks and images separately. When receiving a user query, it retrieves the most relevant content through similarity matching. Finally, the query and sampled tokens are fed into the LLM for answer generation. The detailed procedure is outlined in Algorithm 1 in the appendix.
3.2 Document Parsing
Given a PDF document $D$, the goal of document parsing is to convert it into some machine-readable text blocks $\{T_{1},T_{2},\dots,T_{n}\}$ and diagrams $\{I_{1},I_{2},\dots,I_{m}\}$ according to the reading order and layout structure. By default, text blocks are organized into paragraphs, and all figures and tables are saved as images. These text and images are finally reorganized into an XML file in reading order. This process can be completed using existing open-source PDF parsing tools. During inference, we directly input the parsed structured full data into the subsequent stage of PDF-WuKong.
3.3 Sparse Sampling
For a lengthy multi-page document, if it is directly input into the LLM, there will be two problems. The first is the problem of computing efficiency. The consumption of computing resources will increase dramatically. The second is the problem of inaccurate attention. Key information related to the user query is easily submerged by a large amount of irrelevant content. It is difficult for the model to accurately locate and extract important information in a huge token sequence. Therefore, sparse sampling is essential for efficiently handling lengthy multi-page documents by identifying and extracting the most relevant text chunks or diagrams based on their similarity to the user query.
During the training, for the parsed $n$ text chunks $\{T_{1},T_{2},\ldots,T_{n}\}$, $m$ images $\{I_{1},I_{2},\ldots,I_{m}\}$, and the input user query $q$, we first extract the positive samples and the negative samples for the query. Our PaperPDF dataset has provided corresponding positive single-evidence or multi-evidence samples for each query-answer pair (detailed in Sec. 4). We randomly select two text blocks and two images from the remaining text blocks and images as negative samples. Then, we use a text encoder $En\_T$ to obtain the text embeddings $e_{T_{P}},e_{T_{N}}$ and the query embedding $e_{q}$. An image encoder $En\_I$ is utilized to output the image features $e_{I_{P}},e_{I_{N}}$, which is shared with MLLM.
Given the embeddings of the user query $e_{q}$, the positive samples $E_{P}=\{e_{T_{P}},e_{I_{P}}\}$ and negative samples $E_{N}=\{e_{T_{N}},e_{I_{N}}\}$, we employ a contrastive learning approach to align the text and image features with the query. The goal is to enable the model to capture the document content that is most relevant to the query. The contrastive learning loss is:
$\mathcal{L}_{\text{rep}}=-\frac{1}{P}\sum_{e_{i}\in E_{P}}\log\frac{e^{\frac{\text{sim}(e_{q},e_{i})}{\tau}}}{e^{\frac{\text{sim}(e_{q},e_{i})}{\tau}}+\displaystyle\sum_{e_{j}\in E_{N}}e^{\frac{\text{sim}(e_{q},e_{j})}{\tau}}},$ (1)
where $\text{sim}(e_{q},e_{i})$ and $\text{sim}(e_{q},e_{j})$ represent the similarity between the query and the positive/negative samples. $\tau$ is the temperature parameter that controls the scale of the similarity scores. $P$ is the number of positive samples. By maximizing this probability, the model encourages the representations of the query and positive samples to be closer while pushing the representations of the query and negative samples apart. It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs.
During the inference, we pre-encode all text blocks and images and cache all candidate embeddings. When the user inputs a query, we calculate the similarity between query embedding and cached text/image embeddings. Then the model automatically selects the top-$k$ relevant text blocks and images as evidence to respond to this query. Therefore, this process samples out sparse document content, greatly reducing the computational burden of the subsequent LLM and alleviating the problem of attention shift when facing ultra-long sequences. Moreover, the multimodal embedding cache further optimizes inference time. The pseudocodes for training and inference are shown in the appendix.
3.4 Answer Generation
At this stage, the large language model only receives the document content that is most relevant to the query and discards a lot of redundant information, so it can generate more accurate answers with higher efficiency. Specifically, we input the sampled top-$k$ evidence, the user query, and the task instruction into the LLM, and let it generate an answer based on the provided query and evidence.
Considering that MLLM needs to encode images first for multimodal understanding, we directly input the image tokens obtained from the sparse sampler into the LLM, to save one image encoding process. Thus, the sparse sampler shares the same vision encoder with the MLLM. They can be integrated and trained in an end-to-end manner.
During the training, we input the positive text $T_{P}$ and the positive image tokens $e_{I_{P}}$ into the LLM. Besides, the query and instruction are also input into the LLM. Then, we calculate the cross-entropy loss $\mathcal{L}_{\text{QA}}$ between the output answer $a$ and the ground truth. Finally, the total optimization objective is:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rep}}+\mathcal{L}_{\text{QA}}.$ (2)
PDF-WuKong is optimized end-to-end by these two loss functions for effective multimodal alignment and QA.
4 PaperPDF Dataset
@@
4.1 Overview
Our motivation for creating PaperPDF is threefold: (1) In long document QA contexts, answers often derive from specific segments, with other content acting as noise and complicating MLLM reasoning; (2) Existing datasets are either limited to single-page documents or lack fine-grained evidence ground truth, hampering the training of our sparse sampler; (3) There is currently a lack of a bilingual multimodal PDF QA dataset. Therefore, we introduce a method for automatically generating question-answer pairs from long documents and present PaperPDF, a dataset designed for both training and evaluation.
4.2 Dataset construction strategy
The PaperPDF dataset is constructed through a four-step process as shown in Fig. 3: document parsing, evidence extraction, QA generation, and data filtering. First, $100k$ PDFs are parsed to extract text and image chunks. Then we extract some evidence from each PDF according to our specific rules. Given the evidence and the generation prompt as the input, we use Gemini Pro [43] to generate questions and answers for the training set due to its free and rapid accessibility. GPT-4V [44] is used for the test set to ensure high evaluation quality. After obtaining the preliminary dataset, we conduct automatic filtering and remove abnormal data. For the test data, we further perform manual filtering from multiple aspects to ensure its quality. According to different evidence extraction rules, the triplets in PaperPDF can be divided into two categories.
@@ 4.2 Dataset construction strategy
Single-evidence Q-E-A triplets. The questions in these triplets can be answered based on a single text chunk or a diagram. Therefore, the evidence can be categorized into Text-only and Image-only. These triplets enable PDF-WuKong to initially acquire the capabilities of sparse sampling and long multimodal document understanding.
@@
Multi-evidence Q-E-A triplets. These triplets require reasoning across multiple chunks, involving combinations of text and images. The types include Image-text (derived from a text chunk and associated images), Section (generated from all chunks in a section), and Cross-paragraph (involving related paragraphs across a document). For the third type, semantic summarizations for each paragraph are conducted first, followed by a selection of several related text chunks for QA generation. These multi-evidence triplets enhance our model’s multi-hop reasoning capability.
In total, we obtained $1.1M$ bilingual training data and $10k$ testing data. The dataset statistics are shown in Tab. 2. The appendix contains more data statistics.
@@
Table 2: The statistics of PaperPDF in English and Chinese.
@@ 5 Experiments
5.1 Implementation Details
The LLM and the vision encoder are initialized from IXC2-VL-4KHD [24] and the maximum number of dynamic tiles is set as 16. The text encoder is initialized from BGE-M3 [45]. We train the model by leveraging several document datasets including PaperPDF, DocVQA [35], ChartQA [37], InfoVQA [40], MPDocVQA [46], and DUDE [17]. Before both training and testing, PaperPDF is parsed using Grobid [47] and MinerU [48], while the other datasets are processed following their default instructions. The training is conducted for one epoch using 128 Ascend 910B NPUs with a learning rate of 4e-5. We set the top 5 sampling results as input to the LLM.
@@ 5.1 Implementation Details
IXC2-VL [49] 8B 27.4 30.8 32.8 37.6 21.1 28.5 28.7 30.8 4644
IXC2-VL-RAG [49] 8.5B 32.4 34.0 32.4 48.4 23.3 29.8 32.3 38.1 623
InternVL2 [11] 8B 19.5 28.0 27.6 37.5 24.4 28.2 27.7 30.8 4051
InternVL2-RAG [11] 8B 29.8 29.8 28.3 50.2 24.6 27.8 26.8 43.0 583
@@
Hi-VT5 [46] 0.3B 13.5 3.1 3.7 15.2 - - - - 11589
IXC2-VL [49] 8B 27.1 24.8 25.1 20.5 15.9 19.5 22.2 27.4 4712
InternVL2 [11] 8B 29.4 33.1 35.0 35.5 20.3 33.9 37.8 37.7 5008
@@
Table 3: Performance comparison with open-source models for long PDF understanding on PaperPDF. The best results are marked bold and the second results are underlined.
@@
Gemini pro [50] 26.6 29.0 29.8 67.9
Kimi [51] 28.5 33.6 31.1 74.7
ChatGLM [52] 31.2 35.4 32.0 73.5
Qwen [53] 36.0 40.3 35.5 78.1
@@
Table 4: Performance comparison with commercial products (tested in Sep 2024) for long PDF understanding. The results are tested on a subset of 50 English PDFs. Note: These are the products based on the models rather than the models themselves.
5.2 Long PDF Understanding
To assess the effectiveness of our model in understanding long PDF documents, we conduct comprehensive experiments comparing it with both open-source models and commercial products on the PaperPDF dataset. Due to limited capabilities of traditional document understanding models  [46, 16, 54, 55, 56] and the huge resource costs in advanced MLLMs  [49, 24, 11], achieving deep understanding of lengthy PDF documents remains a highly challenging task. To handle this task, we explore three approaches to input the PDF documents into these MLLMs: pure text content, page images, and parsed interleaved text-image content. For some baselines with plain text input modality, we also report results based on retrieval enhancement techniques  [45]. The experimental results are reported in Tab. 3 and Tab. 4.
Several key conclusions can be drawn from the results. Firstly, inputting parsed interleaved text-image content generally outperforms multiple page images. The main reason is the limited input resolution and number of tokens that the model can accept. Secondly, when handling tokens of similar scale, inputting parsed interleaved text-image content yields better performance compared to pure text content. It is obvious that diagram information in the PDF plays a crucial role for document comprehension. This also indicates that PaperPDF places rigorous requirements on the visual information within documents. Additionally, we observe that for the InternVL2 model, the approach of parsing interleaved text-image content as input outperforms InternVL2-RAG. However, the opposite conclusion is drawn for IXC2-VL. We hypothesize that this discrepancy may be due to the max length setting in IXC2-VL, which could cause the model to overlook some critical information. Finally, benefiting from the inclusion of the spare sampler, our proposed PDF-WuKong model not only surpasses the existing state-of-the-art open-source model InternVL2 by approximately 7% on both the Chinese and English subsets, but also demonstrates competitive performance comparable to proprietary products. Moreover, due to the integration of the sparse sampler, PDF-WuKong maintains efficiency in inference token cost.
@@ 5.2 Long PDF Understanding
Qwen-VL [57] 9.6B 26.4 19.6 18.3
Monkey [58] 9.8B 30.0 24.4 22.3
mPLUG-Owl2 [25] 8.2B 19.5 20.3 22.7
Emu2-Chat [59] 37B 26.0 24.4 23.4
MiniCPM-2.5 [60] 8.5B 31.8 28.2 24.8
IXC2-VL [49] 8B 23.4 20.8 21.3
IXC2-4KHD [24] 8B 24.5 20.0 18.0
CogVLM2 [61] 17B 24.8 27.4 26.3
[…]
To compare with more open-source document MLLMs, considering that most of these models can only handle single-page documents, we construct a subset of the PaperPDF benchmark, only containing test samples with single evidence. Therefore, we provided all models with only one page containing the evidence as their input. The pages are input in the form of images. Our PDF-WuKong can accept input in two formats. One is the page image and another is the parsed page. As shown in Tab. 5, our model’s capability on this subset is significantly better than other document models. Moreover, the document parsing-based paradigm used in our approach is superior to the purely visual paradigm.

### v3

@@ 4.1 Overview
In order to enable PDF-WuKong to simultaneously represent and understand the multimodal content of documents and provide explainable answer evidence, we construct a dataset with English and Chinese academic paper PDFs. The dataset contains complete PDF documents, professional academic questions, answers, and evidence sources for the answers, based on multiple construction strategies. Specifically, our motivation for creating PaperPDF is threefold: (1) In the long document QA task, answers often derive from specific document segments, with other content acting as noise and complicating MLLM reasoning; (2) Existing datasets are either limited to single-page documents or lack fine-grained evidence ground truth, hampering the training of our sparse sampler; (3) There is currently a lack of a bilingual multi-page multimodal document QA dataset. Therefore, we introduce multiple strategies to construct question-answer pairs from long documents for training and evaluation.
@@
4.2 Dataset construction
The academic paper PDF is a kind of typical document that contains rich interleaved text and images, which can intuitively reflect the challenges of our task and the advantages of our model. Therefore, we first obtained approximately 60k English and 10k Chinese documents from open-access academic paper websites, encompassing nearly 70 disciplines, such as computer science, engineering, and materials science.
To build high-quality data, we propose a four-step construction strategy and design a variety of evidence distributions based on human questioning habits. The construction strategy consists of document parsing, evidence extraction, QA generation, and data filtering, as shown in Fig. 3.
First, each PDF is parsed as interleaved text and image chunks according to the statment in Sec. 3.2.
Second, we extract evidence from each PDF according to different extraction rules, and the type of evidence in our PaperPDF can be divided into two categories:
@@ 4.2 Dataset construction
Single-evidence extraction. The evidence can be categorized into Text-only and Image-only. We randomly extract a text paragraph or a diagram from a document and construct questions and answers based only on the information in this single evidence. We stipulate that the length of the extracted paragraphs must contain at least 3 sentences, and the diagram needs to be extracted together with its caption. These enable PDF-WuKong to acquire the initial capabilities of sparse sampling and long multimodal document understanding.
@@
Multi-evidence extraction. The evidence can be categorized into Image-text, Section, and Cross-paragraph. These facilitate reasoning across multiple chunks, involving combinations of text and images. The Image-text evidence is derived from a text chunk and the diagrams it refers to. It is also extracted from a diagram and the text paragraphs that reference it. The Section evidence utilizes all chunks in a section, including all the text paragraphs, figures, and tables. The Cross-paragraph evidence involves related paragraphs across the document. To extract this type of evidence, we first conduct semantic summarizations for each paragraph and then select several of the most semantically relevant text chunks. These multi-evidence samples can significantly enhance our model’s multi-hop reasoning capability.
Third, we leverage commercial MLLM APIs reid2024gemini ; 2023gpt4V to generate questions and answers by feeding the API with extracted evidence and carefully engineered prompts. We implement specific prompts for different types of evidence, including the task definition, requirements and expected output. For question generation, the task definition is to let it generate two academic questions based on the given evidence and require the generation process to rely only on the provided content and eliminate other knowledge. Questions must be brief and objective and fully account for the provided evidence. For answer generation, the model requires thinking step by step based on the question and evidence and generating two answers. The expected outputs include a long answer and a short answer, each with its own thought process. Short answers are used for training and evaluation of PDF-WuKong. Long answers and thought processes can be used to further optimize models to have document reasoning capabilities in the future.
For detailed prompt engineering and the Query-Evidence-Answer (Q-E-A) triplet examples of each type, please refer to the Appendix A. Given the evidence and the prompt as the input, we use Gemini Pro reid2024gemini to generate questions and answers for the training set due to its free and rapid accessibility. GPT-4V 2023gpt4V is used for the test set to ensure high evaluation quality. Using different APIs causes data bias but poses a higher challenge to the generalizability of the model.
Table 2: The statistics of QA pairs for various types of evidence in English and Chinese.
@@
Table 3: Statistical analysis compared with other multi-page document datasets. Page and Token represent the average number of pages per document, and the average number of OCR tokens per document.
MPdataset Doc. QA Lang. Page Token
DUDE van2023document 5k 41.5k En 6 1831
MP-Docvqa TITO2023109834 6k 46k En 8 2026
SlideVQA SlideVQA2023 2.6k 14.5k En 20 1489
PaperPDF (En) 60k 900k En 25 11371
PaperPDF (Zh) 10k 200k Zh 11 3413
PaperPDF 70k 1.1M En+Zh 23 10234
Fourth, to ensure data quality, we conduct data filtering for the Q-E-A triplets. We designed many effective rules to automatically eliminate low-quality samples and remove abnormal data. For example, too-short questions or too-long answers may confuse the model, and duplicate questions will reduce diversity. After performing automatic filtering for the training and testing data, we performed manual filtering and correction for the testing data to improve the quality further. This human verification step is crucial to mitigate potential biases inherent in model-generated content. By having human experts verify and correct the Q-E-A triplets, we ensure that the test set ground truth reflects accurate human reasoning rather than the stylistic patterns of a specific commercial API. We checked if the questions made sense, if the answers were correct, if the Q-E-A triplets were related to each other, etc. We will present the statistical analysis and quality verification of the dataset in the next section.
4.3 Dataset analysis
In total, we obtained $1.1M$ bilingual training data and $10k$ testing data. The number of QA pairs for various types of evidence is shown in Tab. 2. To comprehensively illustrate the characteristics of the PaperPDF dataset, we conduct a detailed statistical analysis in Tab. 3. Compared to previous multi-page document datasets, PaperPDF not only includes a significantly larger number of documents and QA pairs but also features more pages and OCR tokens. More importantly, PaperPDF encompasses both Chinese and English documents with interleaved text and images.
Figure 4: Statistics of question types in PaperPDF dataset.
Fig. 4 shows the distribution of question types in our benchmark. We basically divide the types of questions in multimodal PDF documents into 7 categories, covering various questions that users may encounter. In order to further verify the correctness of the answers corresponding to each type of question, we manually evaluated the answers for each type in the test set. We first randomly selected 100 samples from each type and then asked two doctoral students with scientific research training to give an answer to each question. If the two people’s answers disagree, they need to discuss with each other to determine the final answer. Finally, we counted the consistency between the answers from the test set and the answers given by humans, as shown in Tab. 4. The high agreement in the answers demonstrates the high quality of our benchmark.
Our dataset provides an important innovation driver and a comprehensive benchmark for the development of the multimodal long PDF document understanding task. Explicit constructing Q-E-A triplets aligns with PDF-WuKong’s dual training objectives and ensures PDF-WuKong learns to distinguish fine-grained evidence patterns essential for long-document QA.
Table 4: Answers’ consistency from the testing dataset and human evaluation for different types of questions.
Question Type Consistency
Single-hop understanding 95.2%
Single-hop logical reasoning 93.0%
Multi-hop understanding 93.6%
Multi-hop logical reasoning 92.3%
Figure understanding 94.8%
Table understanding 97.3%
Formula understanding 96.1%
@@ 5 Experiments
Table 5: Performance comparison with open-source models for long PDF understanding on PaperPDF. * indicates that the RAG model has been trained on our dataset. 2p and 3p indicate whether VisRAG retrieves 2 or 3 pages of evidence. The best results are marked bold and the second results are underlined.
@@
IXC2-VL internlmxcomposer2 8B 27.4 30.8 32.8 37.6 21.1 28.5 28.7 30.8 4644
IXC2-VL internlmxcomposer2 +BGE-M3 chen2024bge 8.5B 32.4 34.0 32.4 48.4 23.3 29.8 32.3 38.1 623
InternVL2 chen2024far 8B 19.5 28.0 27.6 37.5 24.4 28.2 27.7 30.8 4051
InternVL2 chen2024far +BGE-M3 chen2024bge 8.5B 29.8 29.8 28.3 50.2 24.6 27.8 26.8 43.0 583
InternVL2 chen2024far +BGE-M3* chen2024bge 8.5B 36.4 37.3 34.7 53.3 28.4 34.2 34.5 46.3 849
@@
Hi-VT5 TITO2023109834 0.3B 13.5 3.1 3.7 15.2 - - - - 11589
IXC2-VL internlmxcomposer2 8B 27.1 24.8 25.1 20.5 15.9 19.5 22.2 27.4 4712
InternVL2 chen2024far 8B 29.4 33.1 35.0 35.5 20.3 33.9 37.8 37.7 5008
Qwen2.5-VL Qwen2.5-VL 7B 35.5 34.8 32.6 38.3 23.5 34.4 38.7 40.8 5878
MiniCPM-V 2.6 yao2024minicpm + VisRAG* yu2024visrag (2p) 10B 31.8 31.0 27.2 27.6 20.5 27.0 30.1 35.9 1444
MiniCPM-V 2.6 yao2024minicpm + VisRAG* yu2024visrag (3p) 10B 32.4 32.5 28.5 28.4 20.8 27.9 30.8 37.7 2199
@@
5.1 Implementation Details
The LLM and the vision encoder are initialized from IXC2-VL-4KHD internlmxcomposer2_4khd and the maximum number of dynamic tiles is set as 16. The text encoder is initialized from BGE-M3 chen2024bge . We train the model by leveraging several document datasets including PaperPDF, DocVQA Mathew_2021_WACV , ChartQA masry2022chartqa , InfoVQA Mathew_2022_WACV , MPDocVQA TITO2023109834 , and DUDE van2023document . Before both training and testing, PaperPDF is parsed using Grobid GROBID and MinerU wang2024mineruopensourcesolutionprecise to support sparse sampling. In contrast, for the other datasets, we utilize the complete full-page images paired with their corresponding OCR text as global input blocks, bypassing the fine-grained parsing step. The training is conducted for one epoch using 128 Ascend 910B NPUs with a learning rate of 4e-5. We set the top 5 sampling results as input to the LLM.
Figure 5: Qualitative comparison between PDF-WuKong with other proprietary products. The red box indicates the evidence that the correct answer depends on. Our model can find the text chunks, figures or tables contaning the evidence and output correct answers.
5.2 Long PDF Understanding
To assess the effectiveness of our model in understanding long PDF documents, we conduct comprehensive experiments comparing it with both open-source models and commercial products on the PaperPDF dataset.
Due to the limited capabilities of traditional document understanding models  TITO2023109834 ; blau2024gram ; huang2022layoutlmv3 ; beltagy2020longformer ; zaheer2020big and the huge resource costs in advanced MLLMs  internlmxcomposer2 ; internlmxcomposer2_4khd ; chen2024far , achieving a deep understanding of lengthy PDF documents remains a highly challenging task. To handle this task, we explore three approaches to input the PDF documents into these MLLMs: pure text content, page images, and parsed interleaved text-image content. For some baselines with plain text modality input, we also report the results based on Text RAG. For the pure vision solutions, we compared some document understanding models that can accept multi-page image input. To ensure a fair comparison regarding data availability, we not only compared against off-the-shelf RAG models but also fine-tuned the retrievers of the RAG baselines (BGE-M3  chen2024bge , VisRAG yu2024visrag ) on the PaperPDF dataset.
Table 6: Comparison with commercial products (tested in Sep 2024) for long PDF understanding. The results are tested on a subset of 50 English PDFs. Note: These are the products rather than the models themselves.
@@ 5.2 Long PDF Understanding
Gemini 1.5 pro gemini 26.6 29.0 29.8 67.9
Kimi kimi4 28.5 33.6 31.1 74.7
ChatGLM chatglm 31.2 35.4 32.0 73.5
Qwen tongyi 36.0 40.3 35.5 78.1
@@
The experimental results for the open-source models are reported in Tab. 5. The performance gap between PDF-WuKong and these fine-tuned baselines cannot be attributed solely to data exposure. Instead, it validates the architectural superiority of our end-to-end sparse sampling mechanism over disjoint retrieval-augmented pipelines. Generally, both document understanding models and RAG models oriented towards a single modality are inferior to our paradigm. Moreover, we compare PDF-WuKong with some closed-source commercial products in Tab. 6, also showing our superiority. The qualitative comparisons between PDF-WuKong and these proprietary products on the long PDF understanding task are presented in Fig. 5 (English) and Fig. 6 (Chinese).
Several key conclusions can be drawn from the results.
1. Interleaved text-image content outperforms plain text input: When handling tokens of similar scale, inputting parsed interleaved text-image content yields better performance compared to pure text content. It is obvious that diagram information in the PDF plays a crucial role in document comprehension. This also indicates that PaperPDF places rigorous requirements on the visual information within documents.
2. Interleaved text-image content outperforms pure vision input: When compared to pure visual input, interleaved text-image content provides superior performance. The main reason is the limited input resolution and number of tokens that the model can handle.
3. Sampling enhances understanding: By introducing the single-modality RAG for MLLM to locate and sample evidence, their ability to understand long documents can be improved, which preliminarily verifies the feasibility of our motivation.
4. Limitations of single-modal sampling: Text-only RAG cannot focus on visual information, while visual-only RAG retrieves page images and cannot extract more fine-grained evidence. The full-page dense text images bring greater challenges to the image encoder.
5. Superior performance on long PDF understanding: Benefiting from the inclusion of the multi-modal spare sampler, our proposed PDF-WuKong not only surpasses InternVL2 by approximately 7% on both the Chinese and English subsets but also demonstrates competitive performance compared with proprietary products. Moreover, due to the integration of the end-to-end sparse sampler, PDF-WuKong maintains efficiency during training and inference with fewer tokens.
@@
Qwen-VL Qwen-VL 9.6B 26.4 19.6 18.3
Monkey li2023monkey 9.8B 30.0 24.4 22.3
mPLUG-Owl2 hu2024mplugdocowl2 8.2B 19.5 20.3 22.7
Emu2-Chat sun2024generative 37B 26.0 24.4 23.4
MiniCPM-2.5 hu2024minicpm 8.5B 31.8 28.2 24.8
IXC2-VL internlmxcomposer2 8B 23.4 20.8 21.3
IXC2-4KHD internlmxcomposer2_4khd 8B 24.5 20.0 18.0
CogVLM2 hong2024cogvlm2 17B 24.8 27.4 26.3
$\text{PDF-WuKong (ours)}^{\dagger}$ 8.5B 36.6 35.2 31.7
$\text{PDF-WuKong (ours)}^{*}$ 8.5B 41.5 42.8 39.8
The main reason for not using our data to train other MLLMs is that they cannot receive lengthy multimodal documents as input. Most models only support single-page input for training and few models can handle up to 8 or 20 pages. They are unable to be trained on our data. To compare with more open-source document MLLMs, we construct a subset of the PaperPDF benchmark, which only contains test samples with single evidence. This subset serves as a critical benchmark for fair comparison. It allows us to compare the fundamental multimodal understanding capabilities of our PDF-WuKong against other models under the same single-page input setting. Therefore, we feed all models with only one page containing the evidence as their input. The pages are input in the form of images. Our PDF-WuKong can accept input in two formats. One is the page image and another is the parsed page. As shown in Tab. 7, our model’s capability on this subset is significantly better than other document models. Moreover, introducing a parsed-input paradigm significantly reduces the difficulty of understanding text-rich documents, yielding approximately a 7% performance improvement on the Single-Evidence subset. This improvement can be primarily attributed to two factors. First, the parsed input allows the model to sample more fine-grained evidence for answer generation. Second, compared to raw full-page images, the parsed image data substantially relaxes the model’s dependency on high-resolution input, thereby enhancing both efficiency and robustness.